# Rendezvous in CAVITY: Kinematics and gas properties of an isolated dwarf-dwarf merging pair in a cosmic void region

### v1

@@ 3 Analysis
To derive the gas and stellar kinematics and measure fluxes of emission lines, we first masked out foreground stars and background galaxies in the FOV of IFU data, and utilized the publicly available full spectral fitting penalized pixel-fitting algorithm (pPXF; Cappellari & Emsellem, 2004; Cappellari, 2017). To estimate the stellar kinematics, pPXF requires a set of sample stellar models for which we used the Vazdekis et al. (2010, 2016) single stellar population (SSP) models based on the MILES stellar library (Sánchez-Blázquez et al., 2006; Cenarro et al., 2007; Falcón-Barroso et al., 2011) with spectral FWHM of 2.51 Å. SSP models used here are based on BASTI isochrones (Pietrinferni et al., 2004) and a bi-modal initial mass function with slope of 1.3 (Vazdekis et al., 1996) and cover the age range of 0.03 to 14.00 Gyr, the metallicity range of -2.27 to 0.40 dex, with the base [/Fe].
@@ 4.1 Kinematics
In the bottom row of the Fig. 3, we show the velocity dispersion of the gas components (i.e., ) and the stellar component (). The maps show fluctuations in values that could be expected in an entangled and complicated interacting system such as here. In particular, regions around both distinct bright central components of this interacting system have the lowest of 705 km s-1 and 15025 km s-1 measured from H and [OIII]5007 emission lines, respectively. Both maps show that toward the system’s outskirts and the faint tail, the values become larger ( 130 km s-1) with average errors of 17 km s-1 and 25 km s-1 for values derived from H and [OIII]5007, respectively. Both maps show the highest values in a small region on the left-hand side, reaching 300 175 km s-1. The detected emission line was best fit with a single Gaussian, and visually, we could not identify any contributions from multiple components in the emission line profiles for any spaxel with S/N ¿ 15 in this system (see Appendix A). Hence, this fluctuation can partly be due to lower S/N in the system’s outskirts. We also believe that the unresolved contributions of both interacting components, which could not be appropriately disentangled with the given data resolution, are partly responsible for larger values in the regions between two bright central components and the outskirts, as they can alter the width of gas emission lines. However, it is important to note that the spectral resolution of the PPAK V500 mode (R 850 at 5000 Å) is significantly lower than the variations and values measured in these maps. Therefore, the results should be interpreted with caution.
We noticed that, through the whole extension of this system, the measured from the [OIII]5007 line (with median of 11618 km s-1) is larger than derived using the H (with median of 827 km s-1). However, this difference (at 2 level) is not significant and might be due to, for example, the presence of multiple interstellar medium (ISM) components heated by different mechanisms (such as star formation) that are unresolved in the present data.
The map is more complicated as it shows higher values in the central regions of component A and much lower values, with a difference of 100 km s-1 in component B. Since CAVITY35843 and CAVITY35844 are in an entangled system with prominent emission lines, pPXF could not retrieve a reliable stellar for most spaxels. Hence, we will not further comment on the measurements.
@@
4.1.1 Separating the kinematics of interacting components
Although components A and B are interacting, their central bright nuclei are still distinct (see Fig. 1). Also, in the kinematic maps of Fig. 3, these two regions seem to be separated with a relative velocity of 100 km s-1 and a fine line of high values that, as discussed before, could be the result of unresolved contributions from both of the interacting components in each spaxel’s spectrum. To take advantage of their unique configuration, we separate the components and investigate their kinematics separately.
@@ 4.1.1 Separating the kinematics of interacting components
In this work, we derive the rotation curves of each component solely from the H emission line (i.e., isolating the H emission line within a 100 Å window around the observed wavelength), as it is the most prominent line in our dataset. Also, based on our results from Fig. 3, we do not expect significant differences in the rotational velocity derived from the Balmer and forbidden lines. Before modelling, we also removed the stellar continuum from the spectra using a first-degree polynomial fit on both sides of the [NII] emission lines444We used specutils.fitting.fit_generic_continuum() function with a linear model to fit the stellar continuum..
@@
In Fig. 4 we present the results of the analysis. On the top panel of this figure, we show the rotational velocity curves for each component. They exhibit a similar range of rotation up to 5 arcsec from the galaxy centre. Beyond this distance, our data only has spatial coverage for component B. Note that, as indicated by the error bars, the uncertainty for values measured beyond this distance is larger, and this is primarily because the analysis was conducted without considering any criteria on spaxels S/N. We do not consider values beyond 5 arcsec in this study, since they are derived based on spaxels with inadequate S/N.
@@
where is the rotation velocity as a function of the radius, , and is the Newtonian constant of gravitation. The curves shown in Fig. 4 indicate that both components have a similar range of dynamical masses up to a galactocentric distance of 5 arcsec, which corresponds to 2.9 kpc. The dynamical mass of each galaxy is, within this range, and , for component A and B respectively. Note that these values represent lower limits to the dynamical mass, as the data used here do not cover the full extent of the emission. Additionally, beam smearing, particularly in cases with low angular resolution (i.e., 2.5 arcseconds in this study) can affect the rotation curves, decreasing their slopes, which may also lower the dynamical masses estimated here. The larger value estimated for the dynamical mass of component A in the very central region may be due to an offset in the estimation of the galaxy’s centre but it is not ultimately reliable physically because the following value has large errorbars. We attempted multiple fits with varying centre estimates, but none resulted in a model that adjusts the galaxy’s rotation better for this region.
The use of for these data is fundamentally risky, as it relies on the assumption that each interacting component exhibits some level of symmetry. Furthermore, the kinematics we are modeling are likely influenced by the conditions under which the two galaxies interacted and the current stage of the merger. While this cannot be thoroughly evaluated with the available data, it undoubtedly contributes to the uncertainty in the results. Additionally, the low resolution of data used here, especially when both galaxies are separated, adds to the uncertainties. Despite these limitations, this test allowed us to confirm the hypothesis of a nearly 1:1 merger within the error bars intrinsic to the nature of the data.
@@
5007 | 2097.7 3.1 | 1273.9 2.9 | 736.6 2.9 |
6583 | 662.4 3.8 | 412.7 3.3 | 207.5 3.4 |
6716 | 510.5 3.1 | 3112.8 3.5 | 172.7 2.8 |
6731 | 357.3 3.1 | 224.8 2.7 | 117.0 2.8 |
Columns are: the name of the emission line, the fluxes measured over the integrated spectra of the entire merging system, as well as for components A and B. All fluxes are corrected for intrinsic reddening.
4.2 Emission lines & Ionisation sources
@@
To facilitate a more accurate comparison with results in the literature, which are primarily based on long-slit spectra of other merging dwarf systems, we constructed integrated spectra for each merging component as well as the entire system. This was achieved by summing the spectral data over spatial elements (spaxels) with S/N 15. For each component, the same masking scheme described in Section 4.1.1 was applied to create the integrated spectra (see also Appendix A). Gas emission fluxes were then derived from these spectra using a strategy similar to that outlined in the beginning of Section 3. The reddening-corrected emission line fluxes obtained from the integrated spectra are listed in Table 2. A comparison of these fluxes, as well as the maps in Fig. 5, particularly for the Balmer lines, clearly indicates that component A is more gas-rich than component B.
@@
To investigate the nature of the ionization mechanisms present in this merging pair we utilised the “Baldwin, Phillips and Terlevich” (BPT) diagnostic diagram (Fig. 6; Baldwin et al., 1981), with the combination of log([OIII]5007/H) vs. log([NII]6583/H). This combination is commonly used to discern the ionisation source of emission lines in galaxies. Emission lines positioned below the Kauffmann et al. (2003)’s empirical division (the black dashed line) are mainly ionised by star formation, and those above the Kewley et al. (2001)’s theoretical one (the gray dashed-dotted line) are AGN dominated. Determining the ionization origin for galaxies (or spaxels, in this case) that fall within the intermediate zone between these two divisions (known as the composite zone) is more complex. Therefore, in the right-hand panel of Fig. 6, we also analysed the combination of log([OIII]5007/H) vs. log([SII]6717,6731/H). As in previous sections, only spaxels with S/N 15 are included in both panels. Most spaxels are located within the star-forming zone, with a few in the AGN zone. The latter spaxels, situated mainly at the outskirts of the interacting system (see Appendix C), are likely affected by low S/N.
@@
Additionally, we measured the equivalent width of the H emission line, EW(H), following Cid Fernandes et al. (2011) to identify any possible weak AGN activity. Spaxels with S/N 15 in this merging system show a range of -0.73 log(EW(H)) 1.46 and -0.81 log([NII]6584/H) -0.53, which confirms that AGN activity is not the ionization source in this system. From Fig. 6 and the EW(H) measurements, it is evident that gas in both components of this system is predominantly ionized by radiation from massive, recently formed stars and possibly by the tidal shocks that are capable of elevating the low-ionization lines such as the [NII] and [SII]. This is further supported by the shallow radio spectral index of -0.630.09 reported by de Gasperin et al. (2018) for this system, which is consistent with a mix of nonthermal radio emission from cosmic-ray electrons, supernova remnants, and thermal bremsstrahlung emission from ionized gas. Thus, we can confidently exclude AGN as a potential ionization source in this system.
@@
Using the stacked spectra of each component, we estimated the stellar mass through full spectral fitting (see Section 3). The derived stellar masses for each component, along with the total stellar mass of the system, are reported in Table 3. These estimates support the findings of Section 4.1.1, confirming that this system is a 1:1 merger. However, it is important to note that these stellar masses are derived from full spectral fitting of the stellar (continuum) component of the merging system. Given that the spectra are heavily dominated by emission lines and the complex, intertwined nature of the data, which makes precise separation of the merging components challenging, these values should be regarded as approximations and interpreted with caution.
@@
The columns include the name of the measured parameter, along with the values obtained for the total system, component A, and component B. The reported dynamical mass is derived within 5 arcseconds of each component. The gas-phase metallicity is based on the empirical calibration of the N2 index (Pettini & Pagel, 2004). The star formation rate (SFR) is calculated using the equation from Kennicutt (1998).
@@ 4.3 Gas metallicity
where N log(). The N empirical calibrator is well suited for the analysis of moderate S/N data and is a better choice for low-mass systems such as CAVITY35843 and CAVITY35844 (e.g., Zurita et al., 2021). A typical systematic uncertainty of 0.2 dex in metallicity derived using the N index is expected according to Pettini & Pagel (2004). In the top panel of Fig. 7 we show the distribution of 12+log(O/H) values measured for each spaxel. We also derived the global gas-phase metallicity from the integrated spectra described in Section 4.2, using the N index. For the integrated spectra of both components, we obtained 12+log(O/H) = 8.55 dex, which matches the value measured from the integrated spectra of component A. Component B shows a slightly lower value of 12+log(O/H) = 8.51 dex (see Table 3). It is evident from the Fig. 7 map and integrated values that, on average, both of the components have similar near-solar gas-phase metallicities, particularly in their central regions, ranging between 8.5012+log(O/H) [dex]8.58, (compared to the solar value of 12+log(O/H) = 8.66 dex; Asplund et al., 2004).
@@
The total stellar mass of this merging system is log(M⋆/M⊙) 9.7 (see Table 1). In a similar mass range, Curti et al. (2020) estimated a value of 12 + log(O/H) = 8.6 dex for star-forming dwarf galaxies using SDSS data. This estimate, like ours, is based on the N empirical calibrator from Pettini & Pagel (2004) and is approximately 0.05 dex higher than the measurements we obtained from the integrated spectrum of this merging pair, as well as each component individually. When employing the O3N2 index from Pettini & Pagel (2004), we find an average 12 + log(O/H) = 8.40 dex for this component, which is still 0.2 dex below Curti et al. (2020)’s findings for star-forming dwarfs in comparable mass ranges using similar calibration methods.
However, these discrepancies do not necessarily imply that the merger event leads to lower gas-phase metallicity in dwarf galaxies, mainly since the 0.05 dex difference we report falls within the typical systematic uncertainty of approximately 0.2 dex associated with the empirical calibrations employed (Pettini & Pagel, 2004). Moreover, further direct comparisons with other interacting or merging dwarf systems are not possible since analysis as such is also limited in the literature and existing studies have employed varied methods for deriving gas-phase metallicity (e.g., Michel-Dansac et al., 2008).
@@ 4.4 Star Formation
Using the extinction corrected H flux, we measure the SFR in this merging system following Kennicutt (1998) equation:
@@
The integrated SFR of this interacting system, whether considered as a whole or for each merging component individually, exceeds the estimates provided by Vilella-Rojo et al. (2021) for star-forming dwarf galaxies with comparable stellar masses, where log(SFR [M⊙ yr-1]), as they reported, is , derived using H emission lines from J-PLUS data (a similar value was also reported using SDSS fibre spectra by Duarte Puertas et al., 2017). Both cited studies rely on one-dimensional spectroscopy and apply methodologies similar to ours in deriving the SFR, which aids in minimizing potential systematic differences when compared with the integrated SFR measured here. These findings are in line with other studies that showed that the interaction between two dwarf galaxies can increase their SFR (e.g., Stierwalt et al., 2015; Gao et al., 2023; Subramanian et al., 2024).
Using the stellar mass estimates, we derived the specific SFR (sSFR = SFR/) of the total system, as well as each merging component (see Table 3). The definition of the star formation main sequence in Janowiecki et al. (2020), places this system in the starburst category, a typical occurrence in gas-rich (wet) mergers (e.g., Hibbard & Mihos, 1995). These findings are also consistent with Lambas et al. (2012) who showed that at a given stellar mass, major mergers are more efficient in forming new stars.
As with the gas-phase metallicity, a direct comparison with the SFR of other merging dwarf systems was not feasible. This is because SFR and sSFR values for many merging/interacting dwarf galaxies are either not reported in the literature or are derived using different methods or indicators (e.g., H vs. UV), rendering meaningful comparisons with our case impractical.

### v2

@@ 3 Analysis
To derive the gas and stellar kinematics and measure fluxes of emission lines, we first masked out foreground stars and background galaxies in the FOV of IFU data, and utilised the publicly available full spectral fitting penalised pixel-fitting algorithm (pPXF; Cappellari & Emsellem, 2004; Cappellari, 2017). To estimate the stellar kinematics, pPXF requires a set of sample stellar models for which we used the Vazdekis et al. (2010, 2016) single stellar population (SSP) models based on the MILES stellar library (Sánchez-Blázquez et al., 2006; Cenarro et al., 2007; Falcón-Barroso et al., 2011) with spectral FWHM of 2.51 Å. SSP models used here are based on BASTI isochrones (Pietrinferni et al., 2004) and a bi-modal initial mass function with slope of 1.3 (Vazdekis et al., 1996) and cover the age range of 0.03 to 14.00 Gyr, the metallicity range of -2.27 to 0.40 dex, with the base [/Fe].
@@ 4.1 Kinematics
In the bottom row of the Fig. 3, we show the velocity dispersion of the gas components (i.e. ) and the stellar component (). The maps show fluctuations in values that could be expected in an entangled and complicated interacting system such as here. In particular, regions around both distinct bright central components of this interacting system have the lowest of 705 km s-1 and 15025 km s-1 measured from H and [OIII]5007 emission lines, respectively. Both maps show that toward the system’s outskirts and the faint tail, the values become larger ( 130 km s-1) with average errors of 17 km s-1 and 25 km s-1 for values derived from H and [OIII]5007, respectively. Both maps show the highest values in a small region on the left-hand side, reaching 300 175 km s-1. The detected emission line was best fit with a single Gaussian, and visually, we could not identify any contributions from multiple components in the emission line profiles for any spaxel with S/N ¿ 15 in this system (see Appendix A). Hence, this fluctuation can partly be due to lower S/N in the system’s outskirts. We also believe that the unresolved contributions of both interacting components, which could not be appropriately disentangled with the given data resolution, are partly responsible for larger values in the regions between two bright central components and the outskirts, as they can alter the width of gas emission lines. However, it is important to note that the spectral resolution of the PPAK V500 mode (R 850 at 5000 Å) is significantly lower than the variations and values measured in these maps. Therefore, the results should be interpreted with caution.
We noticed that through the whole extension of this system, the measured from the [OIII]5007 line (with median of 11618 km s-1) is larger than derived using the H (with median of 827 km s-1). However, this difference (at 2 level) is not significant and might be due to, for example, the presence of multiple interstellar medium components heated by different mechanisms (such as star formation) that are unresolved in the present data.
The map is more complicated as it shows higher values in the central regions of component A and much lower values, with a difference of 100 km s-1 in component B. Since CAVITY35843 and CAVITY35844 are in an entangled system with prominent emission lines, pPXF could not retrieve a reliable stellar for most spaxels. Hence, we do not comment further on the measurements.
@@
4.2 Separating the kinematics of interacting components
Although components A and B are interacting, their central bright nuclei are still distinct (see Fig. 1). Also, in the kinematic maps of Fig. 3, these two regions seem to be separated with a relative velocity of 100 km s-1 and a fine line of high values that, as discussed before, could be the result of unresolved contributions from both of the interacting components in each spaxel’s spectrum. To take advantage of their unique configuration, we separated the components and investigated their kinematics separately.
@@ 4.2 Separating the kinematics of interacting components
In this work, we derived the rotation curves of each component solely from the H emission line (i.e. isolating the H emission line within a 100 Å window around the observed wavelength), as it is the most prominent line in our dataset. Also, based on our results from Fig. 3, we do not expect significant differences in the rotational velocity derived from the Balmer and forbidden lines. Before modelling, we also removed the stellar continuum from the spectra using a first-degree polynomial fit on both sides of the [NII] emission lines444We used specutils.fitting.fit_generic_continuum() function with a linear model to fit the stellar continuum..
@@
In Fig. 4 we present the results of the analysis. On the top panel of this figure, we show the rotational velocity curves for each component. They exhibit a similar range of rotation up to 5 arcsec from the galaxy centre. Beyond this distance, our data only has spatial coverage for component B. We note that, as indicated by the error bars, the uncertainty for values measured beyond this distance is larger, and this is primarily because the analysis was conducted without considering any criteria on spaxels S/N. We do not consider values beyond 5 arcsec in this study, since they are derived based on spaxels with inadequate S/N.
@@
where is the rotation velocity as a function of the radius, , and is the Newtonian constant of gravitation. The curves shown in Fig. 4 indicate that both components have a similar range of dynamical masses up to a galactocentric distance of 5 arcsec, that corresponds to 2.9 kpc. The dynamical mass of each galaxy is, within this range, and , for component A and B respectively. We note that these values represent lower limits to the dynamical mass, as the data used here do not cover the full extent of the emission. Additionally, beam smearing, particularly in cases with low angular resolution (i.e. 2.5 arcseconds in this study) can affect the rotation curves, decreasing their slopes, that may also lower the dynamical masses estimated here. The larger value estimated for the dynamical mass of component A in the very central region may be due to an offset in the estimation of the galaxy’s centre but it is not ultimately reliable physically because the following value has large errorbars. We attempted multiple fits with varying centre estimates, but none resulted in a model that adjusts the galaxy’s rotation better for this region.
The use of for these data is fundamentally risky, as it relies on the assumption that each interacting component exhibits some level of symmetry. Furthermore, the kinematics we are modelling are likely influenced by the conditions under which the two galaxies interacted and the current stage of the merger. While this cannot be thoroughly evaluated with the available data, it undoubtedly contributes to the uncertainty in the results. Additionally, the low resolution of data used here, especially when both galaxies are separated, adds to the uncertainties. Despite these limitations, this test allowed us to confirm the hypothesis of a nearly 1:1 merger within the error bars intrinsic to the nature of the data.
@@
| 5007 | 2097.7 3.1 | 1273.9 2.9 | 736.6 2.9 |
| 6583 | 662.4 3.8 | 412.7 3.3 | 207.5 3.4 |
| 6716 | 510.5 3.1 | 3112.8 3.5 | 172.7 2.8 |
| 6731 | 357.3 3.1 | 224.8 2.7 | 117.0 2.8 |
Measured integrated emission line fluxes. Columns list the name of the emission line, the fluxes measured over the integrated spectra of the entire merging system, and the fluxes for components A and B. All fluxes are corrected for intrinsic reddening.
4.3 Emission lines and ionisation sources
@@
To facilitate a more accurate comparison with results in the literature, which are primarily based on long-slit spectra of other merging dwarf systems, we constructed integrated spectra for each merging component as well as the entire system. This was achieved by summing the spectral data over spatial elements (spaxels) with S/N 15. For each component, the same masking scheme described in Section 4.2 was applied to create the integrated spectra (see also Appendix A). Gas emission fluxes were then derived from these spectra using a strategy similar to that outlined in the beginning of Section 3. The reddening-corrected emission line fluxes obtained from the integrated spectra are listed in Table 2. A comparison of these fluxes, as well as the maps in Fig. 5, particularly for the Balmer lines, clearly indicates that component A is more gas-rich than component B.
@@
To investigate the nature of the ionisation mechanisms present in this merging pair we utilised the ‘Baldwin, Phillips and Terlevich’ (BPT) diagnostic diagram (Fig. 6; Baldwin et al., 1981), with the combination of log([OIII]5007/H) vs. log([NII]6583/H). This combination is commonly used to discern the ionisation source of emission lines in galaxies. Emission lines positioned below the Kauffmann et al. (2003)’s empirical division (the black dashed line) are mainly ionised by star formation, and those above the Kewley et al. (2001)’s theoretical one (the gray dashed-dotted line) are AGN dominated. Determining the ionisation origin for galaxies (or spaxels, in this case) that fall within the intermediate zone between these two divisions (known as the composite zone) is more complex. Therefore, in the right-hand panel of Fig. 6, we also analysed the combination of log([OIII]5007/H) vs. log([SII]6717,6731/H). As in previous sections, only spaxels with S/N 15 are included in both panels. Most spaxels are located within the star-forming zone, with a few in the AGN zone. The latter spaxels, situated mainly at the outskirts of the interacting system (see Appendix C), are likely affected by low S/N.
@@
Additionally, we measured the equivalent width of the H emission line, EW(H), following Cid Fernandes et al. (2011) to identify any possible weak AGN activity. Spaxels with S/N 15 in this merging system show a range of -0.73 log(EW(H)) 1.46 and -0.81 log([NII]6584/H) -0.53, confirming that AGN activity is not the ionisation source in this system. From Fig. 6 and the EW(H) measurements, it is evident that gas in both components of this system is predominantly ionised by radiation from massive, recently formed stars and possibly by the tidal shocks that are capable of elevating the low-ionisation lines such as the [NII] and [SII]. This is further supported by the shallow radio spectral index of -0.630.09 reported by de Gasperin et al. (2018) for this system, which is consistent with a mix of nonthermal radio emission from cosmic-ray electrons, supernova remnants, and thermal bremsstrahlung emission from ionised gas. Thus, we can confidently exclude AGN as a potential ionisation source in this system.
@@
Using the stacked spectra of each component, we estimated the stellar mass through full spectral fitting (see Section 3). The derived stellar masses for each component, along with the total stellar mass of the system, are reported in Table 3. These estimates support the findings of Section 4.2, confirming that this system is a 1:1 merger. However, it is important to note that these stellar masses are derived from full spectral fitting of the stellar (continuum) component of the merging system. Given that the spectra are heavily dominated by emission lines and the complex, intertwined nature of the data, which makes precise separation of the merging components challenging, these values should be regarded as approximations and interpreted with caution.
@@
Physical properties of merging components. The columns include the name of the measured parameter, along with the values obtained for the total system, component A, and component B. The reported dynamical mass is derived within 5 arcseconds of each component. The gas-phase metallicity is based on the empirical calibration of the N2 index (Pettini & Pagel, 2004). The star formation rate (SFR) is calculated using the equation from Kennicutt (1998).
@@ 4.4 Gas metallicity
where N log(). The N empirical calibrator is well suited for the analysis of moderate S/N data and is a better choice for low-mass systems such as CAVITY35843 and CAVITY35844 (e.g. Zurita et al., 2021). A typical systematic uncertainty of 0.2 dex in metallicity derived using the N index is expected according to Pettini & Pagel (2004). In the top panel of Fig. 7 we show the distribution of 12+log(O/H) values measured for each spaxel. We also derived the global gas-phase metallicity from the integrated spectra described in Section 4.3, using the N index. For the integrated spectra of both components, we obtained 12+log(O/H) = 8.55 dex, that matches the value measured from the integrated spectra of component A. Component B shows a slightly lower value of 12+log(O/H) = 8.51 dex (see Table 3). It is evident from the Fig. 7 map and integrated values that, on average, both of the components have similar near-solar gas-phase metallicities, particularly in their central regions, ranging between 8.5012+log(O/H) [dex]8.58, (compared to the solar value of 12+log(O/H) = 8.66 dex; Asplund et al., 2004).
@@
The total stellar mass of this merging system is log(M⋆/M⊙) 9.7 (see Table 1). In a similar mass range, Curti et al. (2020) estimated a value of 12 + log(O/H) = 8.6 dex for star-forming dwarf galaxies using SDSS data. This estimate, similar to ours, is based on the N empirical calibrator from Pettini & Pagel (2004) and is approximately 0.05 dex higher than the measurements we obtained from the integrated spectrum of this merging pair, as well as each component individually. When employing the O3N2 index from Pettini & Pagel (2004), we find an average 12 + log(O/H) = 8.40 dex for this component, which is still 0.2 dex below Curti et al. (2020)’s findings for star-forming dwarfs in comparable mass ranges using similar calibration methods.
However, these discrepancies do not necessarily imply that the merger event leads to lower gas-phase metallicity in dwarf galaxies, mainly since the 0.05 dex difference we report falls within the typical systematic uncertainty of approximately 0.2 dex associated with the empirical calibrations employed (Pettini & Pagel, 2004). Moreover, further direct comparisons with other interacting or merging dwarf systems are not possible since analysis as such is also limited in the literature and existing studies have employed varied methods for deriving gas-phase metallicity (e.g. Michel-Dansac et al., 2008).
@@ 4.5 Star formation
Using the extinction corrected H flux, we measured the SFR in this merging system following the equation of Kennicutt (1998):
@@
The integrated SFR of this interacting system, whether considered as a whole or for each merging component individually, exceeds the estimates provided by Vilella-Rojo et al. (2021) for star-forming dwarf galaxies with comparable stellar masses, where log(SFR [M⊙ yr-1]), as they reported, is , derived using H emission lines from J-PLUS data (a similar value was also reported using SDSS fibre spectra by Duarte Puertas et al., 2017). Both cited studies rely on one-dimensional spectroscopy and apply methodologies similar to ours in deriving the SFR, which aids in minimising potential systematic differences when compared with the integrated SFR measured here. These findings are in line with other studies showing that the interaction between two dwarf galaxies can increase their SFR (e.g. Stierwalt et al., 2015; Gao et al., 2023; Subramanian et al., 2024).
Using the stellar mass estimates, we derived the specific SFR (sSFR = SFR/) of the total system, as well as each merging component (see Table 3). The definition of the star formation main sequence in Janowiecki et al. (2020), places this system in the starburst category, a typical occurrence in gas-rich (wet) mergers (e.g. Hibbard & Mihos, 1995). These findings are also consistent with Lambas et al. (2012) who showed that at a given stellar mass, major mergers are more efficient in forming new stars.
As with the gas-phase metallicity, a direct comparison with the SFR of other merging dwarf systems was not feasible. This is because SFR and sSFR values for many merging and interacting dwarf galaxies are either not reported in the literature or are derived using different methods or indicators (e.g. H vs. UV), rendering meaningful comparisons with our case impractical.